OpenStudy (anonymous):

Is the given ordered pair a solution of the system? (4,10) 9x-y>23 5x+0.2y>20 A. yes B. no

OpenStudy (anonymous):

\[9x-y \ge23\]

OpenStudy (anonymous):

\[5x+0.2y \ge20\]

OpenStudy (igreen):

Plug it in and see if it satisfies the inequalities. (4, 10) \[9x-y \ge23\] \[9(4)-10 \ge23\] Simplify, 9 * 4 = ?

OpenStudy (anonymous):

36

OpenStudy (igreen):

Yes, so we have: \(36 - 10 \ge 23\) WHat's 36 - 10?

OpenStudy (anonymous):

26

OpenStudy (igreen):

Yes, so we have: \(26 \ge 23\) This is true, now let's check the other inequality. \(5x+0.2y\ge20\) \(5(4)+0.2(10)\ge20\) Multiply 5 * 4 and 0.2 * 10 then add them

OpenStudy (anonymous):

20 and 2 so 22

OpenStudy (igreen):

Correct, so we have: \(22 \ge 20\) This is also true, so (4, 10) is a solution to the system of inequalities.

OpenStudy (anonymous):

thanks

OpenStudy (igreen):

No problem.